OpenStudy (anonymous):

Solve the equation sin 5(theta) + sin 3(theta) = 0

OpenStudy (anonymous):

use formula \[\sin(\alpha+\beta)=2\sin(\frac{\alpha+\beta}{2})\cos(\frac{\alpha-\beta}{2})\]

OpenStudy (anonymous):

u will get the answer then

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

you got the answer right??

OpenStudy (anonymous):

I'm on the 3rd step right now but I can't quite close the deal.